OpenStudy (anonymous):

At 2:00pm a car's speedometer reads 50mph, and at 2:10pm it reads 65mph. Use the Mean Value Theorem to find an acceleration the car must achieve.

OpenStudy (jack1):

ok, so do you understand the concept and how Mean Value Theorem works?

OpenStudy (anonymous):

I know that the Mean Value Theorem states that \[\frac{ f(b) -f(a) }{ b-a }=f'(c)\]

OpenStudy (jack1):

yep all it means is that the average rate of change can be approximated to a constant so f(b) in this case is 65 mph f(a) is 50 b is 2:10 a is 2:00 f'(c) is the rate of change parallel to the curve, so in this case: the acceleration b-a = 10 mins = 0.166 hours f(b) - f(a) = 15 15/0.1666 =...?

OpenStudy (anonymous):

Ahhh the only thing I hadn't done when trying to solve it was converting the min to hours lol

OpenStudy (anonymous):

Got it! Thanks so much!

OpenStudy (jack1):

coolski, props ;)